OpenStudy (anonymous):

x^2-10x+25/x-5 simplify

OpenStudy (anonymous):

x - 5

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

\[\frac{x^2-10x+25}{x-5}=\frac{(x-5)(x-5)}{x-5}=x-5\]